# An Entropy-based Text Watermarking Detection Method

### v3

@@ 4.2 Proposed Method
Compute the $z$-statistic:
$\centering z=(|s|_{G}-\gamma\sum_{i=m}^{|T|-1}{W}_{i})/\sqrt{\gamma(1-\gamma)\sum_{i=m}^{|T|-1}{W_{i}}^{2}}\@add@centering$ (4)
if $z>\tau$ then return 1, $i.e.,$Input text is watermarked"
@@
Now we are aware of the positive influence-entropy relation as suggest Eq.3. We decided to take a step further than SWEET: customizing influence weight according to token entropy to fully reflect the influence-entropy relation. While in SWEET, the weight of 1 for above-threshold tokens and the weight of 0 for below-threshold tokens partially reflect the relation, there is not reflection inside the entropy range of above-threshold tokens. Therefore, we utilized a monotonically-increasing and continuous function $f$ to generate weights from token entropy. The weight assigning function can be illustrated as follows:
$W(t)=f(SE(t)-C_{0}).$ (5)
Pseudocode of our proposed detection method is provided in Algorithm 1. We first acquire the text entropy by computing the model logits of each token. The next step is to determine the influence weight of each token, the output of a designed function with their entropy values as inputs. Due to the narrow range of entropy values, our goal here is to create a bigger gap between tokens of high entropy and low entropy. We achieve this by the function $ComputeWeight$, which is training-free and a totally automated process. Here, we first normalize the text entropy. Then we input the processed entropy into a monotonic increasing function, linear for example, to acquire the final weights. After obtaining the entropy-based weights for each token, we then apply the standard detection procedure by Kirchenbauer et al. (2023a) to determine the group of green tokens, $i.e.$, computing the green lists using the detection key and previously generated tokens. Finally, we sum up the weights of the green tokens and calculate the z-statistic using Eq.4 given $|s|_{G}$ and weights of all detected tokens. The detector would give a watermark-positive result if the z-statistic is greater than the given threshold.
In real-world low-entropy scenarios, With our EWD, the vast majority of low-entropy red tokens will not prevent the detector from yielding a positive conclusion. Meanwhile, facing human-written texts, any high-entropy red token will strongly suggest the detector to make a negative output.
@@ 5 Theoretical Analysis
In this section, we theoretically analyzed the Type-I and Type-II error of our proposed detection method. Specifically, we compared EWD with two previous detection methods, KGW and SWEET.
To do so, we adopted Kirchenbauer et al. (2023a)’s steps of approach. Each token in both watermarked and non-watermarked texts can either be green or not, so we can model it as a binomial distribution. While Zhao et al. (2023) set window size in the watermark generator as zero (one universal green list for all tokens) and exhibited major token dependence, $i.e.$, the same tokens in the text are of the same color (either all sampled from the green list or all from the red list), in our case, where a non-zero window size is adopted, the token dependence is significantly weakened. For example, the same token with different prefix would exhibit different sampling result. Therefore, to simplify and facilitate the process of theoretical analysis, we made the reasonable assumption that the sampling of each token is independent. Hence, the sum of green tokens in the text can be approximated with a normal distribution $\mathcal{N}(\mu,\sigma^{2})$ with different sets of mean $\mu$ and variance $\sigma^{2}$ for different detection schemes. Finally, we calculate and compare the probability of the sum surpassing or falling behind the detection threshold $|\tilde{s}|_{G}$, which is the $|s|_{G}$ value in each detection function with a fixed $z$.
@@ 5.1 Type-I Error
In a human-written text, each token is independent of the watermark algorithm. Therefore, its probability to be included in the green list is $\gamma$. Therefore, KGW Kirchenbauer et al. (2023a)’s sum of green tokens is approximated by $\mathcal{N}(\gamma T,\gamma(1-\gamma)T)$. For SWEETLee et al. (2023) that only considers high-entropy tokens, such distribution is modelled as $\mathcal{N}(\gamma\tilde{T},\gamma(1-\gamma)\tilde{T})$, where $\tilde{T}$ is the text length excluding low-entropy tokens. Our method EWD assigned weights to tokens based on entropy, leading to a weighted binomial distribution. We could also approximate it to a normal distribution with mean $\mu=\gamma\sum_{i=m}^{|T|-1}{W}_{i}$ and variance $\sigma^{2}=\gamma(1-\gamma)\sum_{i=m}^{|T|-1}{W_{i}}^{2}$.
When $z$ is set to 2, the probabilities for $|s|_{G}$ under each method to surpass its detection threshold $|\tilde{s}|_{G}$ is identically 2.28%.
@@
5.2 Type-II Error
Type-II error measures the likelihood a machine-generated text is incorrectly identified as human-written. For analysis, we assumed that a sequence generated by the watermark algorithm has a length of $T=200$ tokens and the watermark algorithm has hyper-parameters $\gamma=0.5$ and $\delta=2$.
Originally, Kirchenbauer et al. (2023a) sampled over 500 watermark-generated news scripts and reported an average spike entropy of 0.807. In this section, we focused specifically on cases with a much lower average spike entropy. Based on our data, we approximated the distribution of spike entropy in low-entropy watermarked texts with a power law distribution $f(x)=ax^{a-1}$, where the parameters are $a=0.106$, $loc=0.566$, $scale=0.426$. Under this distribution, the average spike entropy is 0.607. The following probability from Kirchenbauer et al. (2023a) showed a lower bound of the probability of a token $k$ being sampled from the green list:
KGW. Based on Eq.6, the statistical $\mu$ and upper bound of $\sigma^{2}$ can be obtained as $\sum_{k}\mathbb{P}[k\in G]$ and $\sum_{k}\mathbb{P}[k\in G]\cdot(1-\mathbb{P}[k\in G])$. When the average entropy $SE^{*}$ given by the simulated power law distribution is 0.607, we approximated $|s|_{G}$ by $\mathcal{N}(106.9,7.1)$. If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 114) for detection, the false negative rate exceeds 84.6%. This means that KGW is not good at detecting the watermarked texts in cases of low entropy.
SWEET. Following SWEET, we set the entropy threshold at 0.695 and excluded all tokens below this threshold during detection. The average spike entropy $SE^{*}$ and remaining text length $\tilde{T}$ then changed to 0.82 and 24 respectively. Using the equations from KGW, we can get the approximation as $\mathcal{N}(17.5,2.2)$. If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 17.2) for detection, the false negative rate exceeds 42%.
EWD. We assigned higher influence weights to tokens with high spike entropy in order to achieve a better detection rate in cases of low average spike entropy. Note that in experiments, we use several weight functions, and here, for analysis purposes, we used a linear weight function for each token $k$: $W(k)=SE(k)-C_{0}$, where $C_{0}$ is a constant. Based on Eq.6, the statistical $\mu$ and $\sigma^{2}$ can be obtained as $\sum_{k}\mathbb{P}[k\in G]\cdot W(k)$ and $\sum_{k}{W(k)}^{2}\mathbb{P}[k\in G]\cdot(1-\mathbb{P}[k\in G])$. Therefore, we can approximate the $|s|_{G}$ detected by EWD with $\mathcal{N}(5.79,0.56)$. If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 5.55) for detection, the false negative rate exceeds 33.3% (Proof and details in Appendix A). Overall, Our method has a type-II error much lower than KGW and SWEET.
6 Experiments
@@ 6.1 Experiment Settings
Figure 2: The left figure shows the z-scores of watermarked and human texts in the Rotowire, HumanEval and MBPP datasets respectively, each being detected with 3 different methods. The right figure shows the relationship between token weights and the probability of being sampled from the green list in both watermarked and human texts.
@@
Figure 3: We utilized a variety of weight functions and measured their performance in the code detection datasets. The left figure is a visualization of all studied functions, with normalized spike entropy as input. The right figure shows each function’s detection F1 under 1% FPR with comparison to the SWEET baseline. Each data point can correspond to the illustrated function on the left figure by looking at its shape and x-axis value.
Tasks and Datasets. First of all, we used the KGW Kirchenbauer et al. (2023a) to embed watermark in all the following tasks.
@@
To evaluate the detection performance of our proposed method, we targeted at a commonly-seen low-entropy scenario, code generation. We adopted two datasets following Lee et al. (2023), HumanEval Chen et al. (2021) and MBPP Austin et al. (2021). We used StarCoder Li et al. (2023) for code generation. And we used the human-written answers as non-watermarked texts. During evaluation, we restricted the minimal length of both samples to be 15 tokens.
@@
Besides, we fine-tuned a T5-small Raffel et al. (2023) model for data-to-text generation. Specifically, we utilized Wu et al. (2022)’s version of the Rotowire dataset Wiseman et al. (2017), which was converted for a text-to-table task. The dataset consists of NBA game records, presented as tables, and human-written descriptions. We used the table information as input and generate descriptive texts. For each input prompt, the models are expected to generate 200 ± 5 tokens.
We also utilized the pre-trained OPT-1.3B Zhang et al. (2022) for news report generation. Our input prompts are from C4 Raffel et al. (2020) dataset. We restricted the maximum prompt length to 50 tokens. For each input prompt, the models are expected to generate 200 ± 5 tokens using binomial sampling and beam-search.
@@
Baselines and Evaluation Metrics. We selected two detection methods as baselines. The first is KGW detector Kirchenbauer et al. (2023a). The second is SWEET’s detector Lee et al. (2023), which excludes low-entropy tokens during detection. We follow the instructions in SWEET to select entropy threshold from the median, mean, and 3rd quartile (Q3) values of the entropy computed from human texts. For evaluation, we set the false positive rate (FPR) under 5%, following Zhao et al. (2023)’s instruction. We reported the true positive rate (TPR) and F1 score for each method.
@@
Hyper-parameters. For watermarked algorithm, we set $\gamma$ and $\delta$ to 0.5 and 2 respectively. Binomial sampling set the sampling temperature to 0.7. Beam-search set the number of beams to 8. To avoid excessive repetition of text, we set "no_repeated_ngrams" to 16.
@@ 6.2 Main Results
Before detection, we measured the generation quality of all generated texts. Specifically, we are able to achieve 31.0% and 30.4% in pass@1Chen et al. (2021) for HumanEval and MBPP dataset respectively, comparable to Lee et al. (2023)’s. For the Rotowire and C4 dataset, our generated output can achieve 7.24 and 3.25 in text perplexity (PPL) using OPT-6.7B Zhang et al. (2022).
@@
Table 1 shows that for the low-entropy scenario, watermarked code detection, our detector would achieve the best detection accuracy against other baselines. Specifically, for HumanEval and MBPP respectively, our method could achieve a 3.5% and 9.4% improvement than SWEET in true positive rate while maintaining a low-than 5% false positive rate.
@@
Table 1 shows that for the high-entropy scenarios using Rotowire and C4 datasets, our method is overall very similar to other baselines, sometimes even better than them.
6.3 Analysis
@@
Impact of Entropy-based Weights. In figure 2, we gave a few visualizations to show the reason why EWD surpassed other baselines in detection performance. Figure 2(a) shows the z-scores of both watermarked and human texts in different datasets, detected by different methods. Our EWD method would result in overall higher z-score for watermarked texts, and slightly lower z-score for human texts. The z-score increase is more significant in low-entropy code generation tasks than in the Rotowire dataset. By enlarging the gap between the z-scores of watermarked and human texts, our detector could better distinguish the watermarked texts. In figure 2(b), we plotted the relationship of a token’s entropy-based weight and probability of that token being sampled from the green list. In watermarked text, the probability is proportional to the token weight. That is to say, the tokens assigned with a large weight are very likely to be sampled from the green list and contribute to a larger z-score. On the contrary, in human-written texts, there doesn’t exist such a relationship, explaining why the z-scores of human text remain low.
@@
Ablation Study. In the experiments, we utilized a linear function to gain the entropy-based weights for each token. We also studied a variety of monotonic increasing functions as shown in figure 3, which are applied in Eq.5 as $f$. Specifically, each function’s gradient is different. For our custom Sigmoid functions, the gradient starts off high and gradually decreases, that is to say, tokens with relatively low entropy will have significant differences in weights, while tokens with relatively high entropy will exhibit smaller differences. Similarly for our custom Exponential functions, the gradient starts off low and gradually increases, meaning that tokens with relatively low entropy will have smaller differences in weight, while tokens with relatively high entropy will exhibit larger differences. We use function strength to represent how far the function deviates from being linear, for example, "Sigmoid 10" deviates from linear more than "Sigmoid 8".
@@
In the experiments with those functions and two code generation datasets, we can observe that the total 8 listed functions could surpass the SWEET baseline in terms of detection F1. This reflects the versatility of our method, as it is applicable to a wide range of functions.
Time Complexity.
@@
Table 2: The table below shows the average time taken to generate 200-token texts using KGW generator, as well as the average time taken for detection using different methods, measured in seconds (s). The experiment is conducted on a NVIDIA RTX A6000 GPU.
@@
Compared to the KGW detection, both SWEET and our proposed EWD would demand extra time to compute token entropy. Specifically, our detector would also compute weight given the token entropy before calculating the final z-score. According to table 2, the detection time per text is 0.0391, 0.0830 and 0.0831 seconds for KGW, SWEET and our EWD respectively. In spite of being double the value of KGW, our detector remains highly efficient, and the difference with SWEET is virtually insignificant.
Performance against Attacks.
@@
Watermarked texts usually would be edited, which would cause detection performance drop. Here we also evaluated the baseline methods and our EWD under this setting by utilizing back-translation as the attack to remove watermark. Specifically, we firstly generated watermarked texts using the C4 dataset, then translated them from English to French, and later back to English again for detection. In table 3, we can observe that our EWD would achieve the best detection accuracy under 1% FPR and similar results with KGW under 5% FPR.
[…]
Detection w/ General Prompt
Method F1 under 5%FPR Best F1
KGW 0.498 0.741
SWEET 0.580 0.762
EWD 0.600 0.764

### v4

@@ 4.2 Entropy-based Text Watermarking Detection
Compute the detection score $z^{\prime}$ by Eq. (5)
if $z^{\prime}>\tau$ then return 1, $i.e.,$Input text is watermarked"
@@
Pseudocode of our proposed detection method is provided in Algorithm 1. We first acquire the text entropy by computing the model logits of each token. The next step is to determine the influence weight of each token, the output of a designed function ComputeWeight with their entropy values as inputs. After obtaining the entropy-based weights for each token, we then apply the standard detection procedure by Kirchenbauer et al. (2023a) to determine the group of green tokens, $i.e.$, computing the green lists using the detection key and previously generated tokens. Finally, we sum up the weights of the green tokens $|s|_{G}$ and calculate the z-score by
$\centering z^{\prime}=(|s|_{G}-\gamma\sum_{i=m}^{|T|-1}{W}_{i})/\sqrt{\gamma(1-\gamma)\sum_{i=m}^{|T|-1}{W_{i}}^{2}},\@add@centering$ (5)
given $|s|_{G}$ and weight $W_{t}$ of each detected token $t$ in text $T$. The detector would give a positive result if the z-score is greater than the given threshold. It is noticably that the proposed algorithm is a fully automated process that eliminates the need for manual operations from the SWEET method.
@@ 5 Theoretical Analysis
In this section, we theoretically analyze the Type-I and Type-II error of our proposed detection method. Specifically, we compare EWD with two previous detection methods, KGW and SWEET.
To do so, we adopt the same steps as in KGW Kirchenbauer et al. (2023a). Each token in both watermarked and non-watermarked texts can either be green or red, so we can model it as a binomial distribution. While Zhao et al. (2023) set the window size in the watermark generator as zero (one universal green list for all tokens) and exhibits major token dependence, $i.e.$, the same tokens in the text are of the same color, in our case, where a non-zero window size is adopted, the token dependence is significantly weakened. For example, the same token with different prefix would exhibit different sampling result.
Therefore, to simplify and facilitate the process of theoretical analysis, we make the reasonable assumption that the sampling of each token is independent. Hence, the sum of green tokens in the text can be approximated with a normal distribution $\mathcal{N}(\mu,\sigma^{2})$ with different sets of mean $\mu$ and variance $\sigma^{2}$ for different detection schemes. Finally, we calculate and compare the probability of the sum surpassing or falling behind the detection threshold $|\tilde{s}|_{G}$, which is the $|s|_{G}$ value in each detection function with a fixed $z$.
@@ 5.1 Type-I Error
In a human-written text $T$, each token is independent of the watermark algorithm. Therefore, its probability to be included in the green list is $\gamma$. Therefore, KGW’s sum of green tokens is approximated by $\mathcal{N}(\gamma|T|,\gamma(1-\gamma)|T|)$. For SWEET that only considers high-entropy tokens, such distribution is modelled as $\mathcal{N}(\gamma\tilde{|T|},\gamma(1-\gamma)\tilde{|T|})$, where $\tilde{T}$ is the text excluding low-entropy tokens. Our method EWD assign weights to tokens based on entropy, leading to a weighted binomial distribution. We could also approximate it to a normal distribution with mean $\mu=\gamma\sum_{i=m}^{|T|-1}{W}_{i}$ and variance $\sigma^{2}=\gamma(1-\gamma)\sum_{i=m}^{|T|-1}{W_{i}}^{2}$.
The theoretical Type-I errors for all detection methods are the same. When the detection threshold is set to 2, the probabilities for $|s|_{G}$ under each method to surpass its detection threshold $|\tilde{s}|_{G}$ is identically 2.28%.
5.2 Type-II Error
Type-II error measures the likelihood a machine-generated text is incorrectly identified as human-written. For analysis, we assume that a sequence generated by the watermark algorithm has a length of $|T|=200$ tokens and the watermark algorithm has hyper-parameters $\gamma=0.5$ and $\delta=2$.
Originally, KGW samples over 500 watermark-generated news scripts and reports an average spike entropy of 0.807. In this section, we focus specifically on cases with a much lower average spike entropy. Based on our data, we approximate the distribution of spike entropy in low-entropy watermarked texts with a power law distribution $f(x)=ax^{a-1}$, where the parameters are $a=0.106$, $loc=0.566$, $scale=0.426$. Under this distribution, the average spike entropy is 0.608. The following probability from KGW shows a lower bound of the probability of a token $k$ being sampled from the green list:
KGW. Based on Eq. (6), the statistical $\mu$ and upper bound of $\sigma^{2}$ can be obtained as $\sum_{k}\mathbb{P}[k\in G]$ and $\sum_{k}\mathbb{P}[k\in G]\cdot(1-\mathbb{P}[k\in G])$. When the average entropy $SE^{*}$ given by the simulated power law distribution is 0.608, we approximate $|s|_{G}$ by $\mathcal{N}(107.10,49.70)$. If we set a threshold of $z$ = 2 (corresponding to $|\tilde{s}|_{G}$ = 114.14) for detection, the false negative rate exceeds 84.1%. This means that KGW is not good at detecting the watermarked texts in cases of low entropy.
SWEET. Following SWEET, we set the entropy threshold at 0.695 and exclude all tokens below this threshold during detection. The average spike entropy $SE^{*}$ and remaining text length $\tilde{|T|}$ then change to 0.82 and 24, respectively. Using the equations from KGW, we can get the approximation as $\mathcal{N}(17.48,4.84)$. If we set a threshold of $z$ = 2 (corresponding to $|\tilde{s}|_{G}$ = 17.02) for detection, the false negative rate exceeds 41.7%.
EWD. EWD assigns higher influence weights to tokens with high spike entropy. Note that in experiments, we use several weight functions, and here, for analysis purposes, we use a linear weight function for each token $k$: $W(k)=SE(k)-C_{0}$, where $C_{0}$ is a constant. Based on Eq. (6), the statistical $\mu$ and $\sigma^{2}$ can be obtained as $\sum_{k}\mathbb{P}[k\in G]\cdot W(k)$ and $\sum_{k}{W(k)}^{2}\mathbb{P}[k\in G]\cdot(1-\mathbb{P}[k\in G])$. Therefore, we can approximate the $|s|_{G}$ detected by EWD with $\mathcal{N}(5.79,0.31)$. If we set a threshold of $z^{\prime}$ = 2 (corresponding to $|\tilde{s}|_{G}$ = 5.55) for detection, the false negative rate exceeds 33.4% (details in Appendix A).
Table 2 summarizes our calculated theoretical Type-I and Type-II error using different methods in low-entropy detection scenarios. Overall, our method achieves the lowest Type-II error among all methods while maintaining the same Type-I error.
Table 2: This table summarizes our calculated theoretical Type-I and Type-II error using different methods in low-entropy detection scenarios. The data in the table is presented in percentages (%). Lower values indicates better detection accuracy.
Methods Type-I Error Type-II Error
KGW 2.28 84.1
SWEET 2.28 41.7
EWD 2.28 33.4
6 Experiments
@@
In this section, we first introduce the tasks and datasets used for watermarked text generation. Then, we compare the performance of our EWD against several baselines on those datasets. Next, we analyze the role of our proposed entropy-based weight and study the performance of EWD using different weight computing functions. Finally, we evaluate the detection performance without the original prompt, detection speed and robustness of our EWD.
@@ 6.1 Experiment Settings
Figure 2: Subfigure (a) shows the z-scores of watermarked and human texts in the Rotowire, HumanEval and MBPP datasets, respectively, each being detected with 3 different methods. Subfigure (b) shows the relationship between token weights and the probability of being green in both watermarked and human texts.
@@
Figure 3: We utilize two additional weight functions other than Linear and measured their performance in the code detection datasets. Subfigure (a) is a visualization of all studied functions, with normalized spike entropy as input. Subfigure (b) shows each function’s detection F1 under 1% FPR with comparison to the SWEET baseline. Each data point can correspond to the illustrated function on the left figure by looking at its shape and x-axis value.
Tasks and Datasets. There are various text generation tasks in the field of Natural Language Processing, and here we aim to generate watermarked text of both low-entropy and high-entropy distributions. Since our focus is on comparing different detection methods, we utilize the same KGW watermark generation algorithms in all the following tasks and datasets.
@@
For low-entropy scenarios, we target at the task of code generation by adopting two datasets following Lee et al. (2023), HumanEval Chen et al. (2021) and MBPP Austin et al. (2021). The two datasets contain test cases of python programming problems and corresponding reference answers, which are utilized as human-written samples. Code generation is conducted by StarCoder Li et al. (2023). During evaluation, the length of both samples are restricted to be at least 15 tokens.
@@
The first common high-entropy scenario is the news report generation. Our input prompts to the pre-trained OPT-1.3B Zhang et al. (2022) are from the C4 Raffel et al. (2020) dataset. Besides, we include another high-entropy data-to-text generation task by utilizing Wu et al. (2021)’s version of the Rotowire dataset Wiseman et al. (2017). The dataset consists of NBA game records, presented as tables, and human-written descriptions. We fine-tune a T5-small Raffel et al. (2023) model for text generation and input the table information to the fine-tuned model. We generate 200 samples of length 200 ± 5 tokens for each high-entropy tasks, using binomial sampling and beam search.
@@
Baselines and Evaluation Metrics. Two detection methods are selected as baselines. The first is KGW detector. The second is SWEET’s detector, which excludes low-entropy tokens during detection. Note that SWEET proposes both selective watermark generation and detection, while here we only adopt the detection part. The selection of the entropy threshold follows the instructions in SWEET. For evaluation, we set the false positive rate (FPR) under 5%, following Zhao et al. (2023)’s instruction. We report the true positive rate (TPR) and F1 score for each method.
@@
Hyper-parameters. For the KGW generator, $\gamma$ and $\delta$ is set to 0.5 and 2, respectively. Binomial sampling sets the sampling temperature to a common value of 0.7. Beam-search sampling sets the number of beams to 8. To avoid excessive repetition of text, "no_repeated_ngrams" is set to 16.
@@ 6.2 Main Results
Before detection, we measure the generation quality of all generated texts. Specifically, we are able to achieve 31.0% and 30.4% in pass@1Chen et al. (2021) for the HumanEval and MBPP dataset, respectively, comparable to Lee et al. (2023)’s. For the Rotowire and C4 dataset, our generated output can achieve 7.24 and 3.25 in text perplexity (PPL) using OPT-6.7B Zhang et al. (2022).
@@
Table 3 shows that for the low-entropy scenario, watermarked code detection, our detector would achieve the best detection accuracy against other baselines. Specifically, for HumanEval and MBPP, respectively, our method could achieve a 2.5% and 9.4% improvement than SWEET in TPR while maintaining a low-than 5% FPR.
@@
For the high-entropy scenarios using the Rotowire and C4 datasets, the detection performance of our method is overall very similar to other baselines, sometimes even better than some baselines.
6.3 Empirical Analysis
@@
Impact of Entropy-based Weights. In Figure 2, we provide a few visualizations to show the reason why EWD surpasses other baselines in detection performance. Figure 2(a) shows the z-scores of both watermarked and human texts in different datasets, detected by different methods. Our EWD method would result in overall higher z-scores for watermarked texts, and slightly lower z-scores for human texts. The z-score increase is more significant in low-entropy code generation tasks than in the Rotowire dataset. By enlarging the gap between the z-scores of watermarked and human texts, our detector could better distinguish the watermarked texts. In Figure 2(b), we plot the relationship of a token’s entropy-based weight and probability of that token being sampled from the green list. In human-written texts, there does not exist a relationship between the probability and token weights, explaining why the z-scores of human text remain low. Meanwhile, in watermarked texts, the probability is proportional to the token weight, which means the tokens assigned with a large weight are very likely to be sampled from the green list and contribute to a larger z-score. This proves that our entropy-based weights can help more accurately reflect the watermark level in low-entropy texts.
@@
Performance with Different Weight Functions. In the experiments, we utilize a linear function to gain the entropy-based weights for each token. We also study two more monotonic increasing functions, the Sigmoid and the Exponential function. As shown in Figure 3(a), we further change the function strength to 4 different values in each function, which represents how far the function deviates from being linear, for example, "Sigmoid 10" deviates from linear more than "Sigmoid 8". The abovementioned functions are applied in Eq. (4) as $f$. Specifically, each function’s gradient is different. For our custom Sigmoid functions, the gradient starts off high and gradually decreases, meaning that tokens with relatively low entropy will have significant differences in weights, while tokens with relatively high entropy will exhibit smaller differences. Similarly, for our custom Exponential functions, the gradient starts off low and gradually increases, meaning that tokens with relatively low entropy will have smaller differences in weight, while tokens with relatively high entropy will exhibit larger differences.
@@
Figure 3(b) demonstrates that the two listed functions with different strength could surpass the SWEET baseline in terms of detection F1. This reflects the versatility of our method, as it is applicable to more functions.
Detection w/ General Prompt
Methods F1 under 5%FPR Best F1
KGW 0.498 0.741
SWEET 0.580 0.762
EWD 0.600 0.764
[…]
Detection Speed.
@@
Table 5: This table shows the average time taken to generate 200-token texts using KGW generator, as well as the average time taken for detection using different methods, measured in seconds. The experiment is conducted on a NVIDIA RTX A6000 GPU.
@@
Compared to the KGW detection, both SWEET and our proposed EWD would demand extra time to compute token entropy. Specifically, our detector would also compute weights given the token entropy before calculating the final z-score. According to Table 5, the detection time per text is 0.0391, 0.0830 and 0.0831 seconds for KGW, SWEET and our EWD, respectively. In spite of being double the value of KGW, our detector remains highly efficient, and the difference with SWEET is virtually insignificant.
Performance against the Back-translation Attack.
@@
Watermarked texts usually would be edited before being detected, which would remove part of the watermark and cause detection performance drop. Here we also evaluate the baseline methods and our EWD under this setting by utilizing back-translation as the attack to remove watermark. Specifically, we firstly generate watermarked texts using the C4 dataset, then translate them from English to French, and later back to English again for detection. In Table 6, our EWD would achieve the best detection accuracy under 1% FPR and similar results with KGW under 5% FPR.